Three numbers are positive consecutive odd integers. If the larger two numbers have a product of 195, find all three integers.
Answer by texttutoring(324)

Let x = the first integer
Let x + 2 = the second integer (because 3 is 2 more than 1, and 5 is 2 more than 3, and so on...)

Let x + 4 = the 3rd integer

The two larger numbers (x+2) and x+4) multiply to 195.
195 = (x+2)(x+4)
195 = x^2 + 6x + 8

You have to factor this to solve it:
0 = x^2 +6x +8 -195
0 = x^2+6x-187

To factor this, you have to find 2 numbers that multiply to -195 and have a difference of 6. Or you could use the Quadratic Equation, with a=1 b=6 and c=-187

Either way you will find that those numbers are -17 and 11.
The three numbers are either 11, 13, and 15 or -17, -15, and -13.